Use Coulomb's law to determine the magnitude of th
Answer #1

At points A, we have

d2 = (2a)2 + a2 = (0.094 m)2 + (0.047 m)2

d2 = 0.011045 m2

using a trigonometric identity, we have

\theta = tan-1 [(0.047 m) / (0.094 m)]

\theta = 26.5 degree

using a formula, we have

EA = ke Q / d2 = (9 x 109 Nm2/C2) (6.1 x 10-6 C) / (0.011045 m2)

EA = 4970.5 x 103 N/C

EA = 4.97 x 106 N/C

The x-component of an electric fied is given by,

EA,x = (4.97 x 106 N/C) cos 26.50 = 4.44 x 106 N/C

The y-component of an electric fied is given by,

EA,y = (4.97 x 106 N/C) sin 26.50 = 2.21 x 106 N/C

Since the magnitude left will eventually cancel out with the magnitude coming from the Charge on other side of A, we can ignore that number.

The Charge on left side will yield the same magnitude up. So at this point, we can just multiply by two to get the answer -

E = 2.21 x 106 N/C

At points B, we have

For right side ; r2 = (3a)2 + a2 = [(0.141 m)2 + (0.047 m)2]

r2 = 0.02209 m2

using a trigonometric identity, we have

\theta = tan-1 [(0.047 m) / (0.141 m)]

\theta = 18.4 degree

We know that,   EB = ke Q / r2 = (9 x 109 Nm2/C2) (6.1 x 10-6 C) / (0.02209 m2)

EB = 2485.2 x 103 N/C

EB = 2.48 x 106 N/C

The x-component of an electric fied is given by,

EB,x = (2.48 x 106 N/C) cos 18.40 = 2.35 x 106 N/C      (left)

The y-component of an electric fied is given by,

EB,y = (2.48 x 106 N/C) sin 18.40 = 0.782 x 106 N/C      (up)

For left side ;    r2 = [(0.047 m)2 + (0.047 m)2]

r2 = 0.094 m2

using a trigonometric identity, we have

\theta = tan-1 [(0.047 m) / (0.047 m)]

\theta = 45 degree

We know that,   EB = ke Q / r2 = (9 x 109 Nm2/C2) (6.1 x 10-6 C) / (0.094 m2)

EB = 584.04 x 103 N/C

EB = 5.84 x 105 N/C

The x-component of an electric fied is given by,

EB,x = (5.84 x 105 N/C) cos 450 = 4.12 x 105 N/C       (right)

The y-component of an electric fied is given by,

EB,y = (5.84 x 105 N/C) sin 450 = 4.12 x 105 N/C      (up)

To sum the y-component of an electric field -

EB,y = [(0.782 x 106 N/C) + (4.12 x 105 N/C) = 1194000 N/C

To sum the x-component of an electric field -

EB,x = [(2.35 x 106 N/C) + (-4.12 x 105 N/C)]

EB,x = 1938000 N/C

Magnitude :    EB = \sqrt{}(1938000 N/C)2 + (1194000 N/C)2

EB = 2.27 x 106 N/C

Using a trigonometric identity, we have

\theta = tan-1 [(1194000 N/C) / (1938000 N/C)]

\theta = 31.6 degree